OpenStudy (anonymous):

Given the equation \[-4 \sqrt{x-3} = 12\] , solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@campbell_st

OpenStudy (jessiegonzales):

is there an answer choice as no solution

OpenStudy (anonymous):

x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (aum):

Square both sides. What do you get?

OpenStudy (anonymous):

Those are the possible answers

OpenStudy (anonymous):

-4 (x-3) = 144???

OpenStudy (aum):

Yes. Distribute the -4. What do you get?

OpenStudy (anonymous):

-4x + 12 = 144

OpenStudy (anonymous):

Oh I understand it now thank you!!!!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

This is wrong

OpenStudy (anonymous):

You need to divide by -4 before you square both sides

OpenStudy (aum):

Oh, it should be 16 not 4 when you squared.

OpenStudy (aum):

Square both sides: 16(x - 3) = 144 divide both sides by 16: x - 3 = 9 x = 12

OpenStudy (anonymous):

Or do that, but dividing it by -4 would be much simpler

OpenStudy (aum):

Put 12 back in the original equation: -4 * 3 = -12 which is not equal to 12. So x = 12 is an extraneous solution.

OpenStudy (anonymous):

Aum, show your work for that last one, please? I can't see how you got there